has an x-intercept of 2 and is perpendicular to the graph of 5x-3y=2
Answer by venugopalramana(3286)

5x-3y=2..perpendicular line to this is 3x+5y=k..its x intercept is 2 hence if y=0 then x=2
3*2+5*0=k=6..eqn of the reqd.line is 3x+5y=6